OpenStudy (anonymous):

Find the integral of (3^(1/2) +x^(1/2))^2

OpenStudy (mimi_x3):

Is it: \[\large \int\limits(\sqrt{3} +\sqrt{x} )^{2} \]??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[ \int(x+2\sqrt{3x}+3)dx=\int x\ dx+2\sqrt3\int\sqrt x\ dx+3\int dx \]

OpenStudy (anonymous):

could explain some of the steps that you took?

OpenStudy (anonymous):

\[ (\sqrt3+\sqrt x)^2=(\sqrt3+\sqrt x)(\sqrt3+\sqrt x)=x+2\sqrt{3x}+3 \]

OpenStudy (anonymous):

oooh okay i see so you expanded it first.. alright i think i get this question now. Thanks !